600.0 mL of air is at 20.0 °C. What is the volume at 60.0 °C?
Answer #1

According to Charles's Law, the equation is as

V1/T1 = V2/T2

V1 = 600.0 mL ; V2 = ?

T1 = 273 + 20 = 293 K

T2 = 273 + 60 = 333 K

(600.0 mL)/(293 K) = V2/(333 K)

V2 = (333 K)(600.0 mL)/(293 K) = 681.9 mL